OpenStudy (anonymous):

If \(\gcd(a,30)=1\), show that \(60\) divides \(a^4+59\)

OpenStudy (anonymous):

60 = 2^2 . 3 . 5

OpenStudy (anonymous):

\(\large a\) is not even, so \(\large a \equiv 1 \pmod 2 \implies a^2 \equiv 1 \pmod 4 \implies a^4 \equiv 1 \pmod 4\)

OpenStudy (ikram002p):

Hint :- a is odd thus a=1 mod 4 or a=-1 mod 4 both cases gives a^4=1 mod 4 now a^4=1 mod 4 (Fermat ) a^2=1 mod 3 a^4=1 mod 3

OpenStudy (anonymous):

got it, thanks!

OpenStudy (ikram002p):

not welcome

OpenStudy (ikram002p):

i made a typo with next step a^4=1 mod 5

OpenStudy (anonymous):

How is below argument without using odd/even thingy : \(\large a = 1 \pmod 2\\ \implies a^2 = 1 \pmod 2 \\ \implies a^2-1 \equiv 0 \pmod 2 \\ \implies (a+1)(a-1)\equiv 0 \pmod 2\)

OpenStudy (ikram002p):

a is odd , so a-1 and a+1 both are even

OpenStudy (anonymous):

Exactly! so both are divisible by 2 and hence (a+1)(a-1) = 0 mod 2^2

OpenStudy (ikram002p):

yep.